# Decentralized Optimization with Coupled Constraints

### v3

@@ B.5 Lower bound on the number of gradient computations
$\min_{x_{1},\ldots x_{n}\in\mathbb{R}^{d}}~{}$ $\sum_{i=1}^{n}f_{i}(x_{i})+\sum_{i=1}^{n}v_{i}(u_{i})$
s.t. $\sum_{i=1}^{n}A_{i}x_{i}=0$
where all $v_{i}(u)$ are the same and $v_{i}(u_{i})=g(u)=\sum_{j=1}^{n}g_{j}(u)$ and $g_{j}(u)$ are defined in (39). All $v_{i}(u)$ are the same and are defined as
$v_{i}(u_{i})=\frac{1}{2}u_{i}^{\top}\mathbf{M}u_{i}+\frac{\mu_{f}}{L_{f}}u^{\top}u-u_{1}.$
Since there is no communication constraint on $u_{i}$, each node runs optimization process individually. Following the same arguments as for function $g(z)$, we get the lower bound on the number of oracle calls

### v4

@@ C.5 Lower bound on the number of gradient computations
$\begin{aligned} \min_{\begin{subarray}{c}x_{1},\ldots x_{n}\in\mathbb{R}^{d}\\ u_{1},\ldots,u_{n}\in\mathbb{R}^{d}\end{subarray}}~&\sum_{i=1}^{n}f_{i}(x_{i})+\sum_{i=1}^{n}v_{i}(u_{i})\\ \text{s.t. }~&\sum_{i=1}^{n}\mathbf{A}_{i}x_{i}=0\end{aligned}$ (47)
where all $f_{i}(x)$ are defined in (39) and all $v_{i}(u_{i})$ are similar and defined as
$v_{i}(u_{i})=\frac{L_{f}}{8}u_{i}^{\top}\mathbf{M}u_{i}+\frac{\mu_{f}}{2}u^{\top}u-\frac{L_{f}}{4}u_{1}.$
First, note that each $v_{i}(u_{i})$ is $L_{f}$-smooth and $\mu_{f}$-strongly convex, i.e. problem (47) satisfies the assumptions of Theorem 2.
Problem (47) falls into two independent parts. The first part is minimization of $\sum_{i=1}^{n}f_{i}(x_{i})$ subject to constraints and is identical to (35), while the second part is minimization of $\sum_{i=1}^{n}v_{i}(u_{i})$ without constraints. Therefore, the lower bounds on the number of communications and number of matrix multiplications for problem (47) are inherited from lower bounds for problem (35) and are the given by $N_{\mathbf{W}}$ in (45) and $N_{\mathbf{A}}$ in (46). It remains to lower bound the number of oracle calls for minimization of $\sum_{i=1}^{n}v_{i}(u_{i})$.
The structure of $v_{i}(u_{i})$ is the same as the structure of $\sum_{j=1}^{n}g_{j}(z)$ defined in (41). Therefore, the minimum of each $v_{i}(u_{i})$ is attained at the same point $u_{i}^{*}=u^{*}$, and the $k$-th component of $u^{*}$ is given by $(u^{*})_{k}=\nu^{k}$, where
$\nu=\frac{\sqrt{\frac{L_{f}}{\mu_{f}}+\frac{2}{3}}-1}{\sqrt{\frac{L_{f}}{\mu_{f}}+\frac{2}{3}}+1}.$
Since there is no communication constraint on $u_{i}$, each node runs optimization process individually. Following the same arguments as for function $\sum_{j=1}^{n}g_{j}(z)$, we get the lower bound on the number of oracle calls